OpenStudy (anonymous):

Help pls What is the line of symmetry for the parabola whose equation is y = -x^2 + x + 3? The line of symmetry for the quadratic equation y = ax^2 + 8x - 3 is x = 4. What is the value of "a"? Find the vertex of the parabola whose equation is y = x^2 + 6x + 2. What is the line of symmetry for the parabola whose equation is y = 3x^2 + 24x - 1? What is the line of symmetry for the parabola whose equation is y = x^2 + 10x + 25? Find the vertex of the parabola whose equation is y = -2x^2 + 8x - 5. What is the line of symmetry for the parabola whose equation is y = x^2

OpenStudy (anonymous):

whoa nellie how many questions you got? lets do the first one

OpenStudy (anonymous):

this is 7 out of 10 that i got wrong on a test i have for an online class :'[

OpenStudy (anonymous):

line of symmetry of \(y=ax^2+bx+c\) is \(x=-\frac{b}{2a}\) line of symmetry for \(y=-x^2+x+3\) is \(x=\frac{-1}{2\times (-1)}=\frac{1}{2}\)

OpenStudy (anonymous):

first one is therefore \(x=\frac{1}{2}\)

OpenStudy (anonymous):

thank you @satellite73

OpenStudy (anonymous):

want to try the next one?

OpenStudy (anonymous):

i dont know what to do with a and x

OpenStudy (anonymous):

\[ y = ax^2 + 8x - 3 \] you don't do anything with the \(x\) you are told that the axis of symmetry is \(x=4\) making \(-\frac{b}{2a}=4\)

OpenStudy (anonymous):

since \(b=8\) you have \[-\frac{8}{2a}=4\] and you can solve that equation for \(a\)

OpenStudy (anonymous):

there isnt a b there..

OpenStudy (anonymous):

whoa lets back up general form is \[\large y=\color{red}ax^2+\color{blue}bx+\color{green}c\] you have \[\large y=\color{red}ax^2+\color{blue}8x+\color{green}{-3}\]

OpenStudy (anonymous):

that makes \(\color{blue}b=\color{blue}8\)

OpenStudy (anonymous):

oh i see

OpenStudy (anonymous):

is a 1?

OpenStudy (anonymous):

now your job is to solve \[-\frac{8}{2a}=4\] for \(a\)

OpenStudy (anonymous):

not quite don't forget the \(-\) sign

OpenStudy (anonymous):

no wait i got it a is -2

OpenStudy (anonymous):

-1*

OpenStudy (anonymous):

\[-\frac{8}{2a}=4\\ -8=8a\\ -1=a\]

OpenStudy (anonymous):

right

OpenStudy (anonymous):

i meant to put 1

OpenStudy (anonymous):

k good

OpenStudy (anonymous):

thanks! @satellite73

OpenStudy (anonymous):

there is this really cool program called geogebra, you can type equations in and it will graph, it helped me a lot with these types of problems,

OpenStudy (anonymous):

post the next one in a new thread and i will help you with it